At the beginning of each month an amount of X rand is deposited into a savings account earning k × 100% interest per year, compounded monthly. The future value of these savings after 24 deposits can be denoted by

[1] S = X(1 + k 12 ) 24 .

[2] S = (1 + k 12 )Xs 24 k÷12.

[3] S = Xs 24 k .

[4] S = (1 + k)Xs 24 k .

[5] none of the above


1
Expert's answer
2021-06-06T15:59:17-0400

[5] none of the above


This is a future annuity due which will be given by


Future Value of an Annuity


Due (FVAD) Formula"=A×\frac{(1 + r)^n - 1}{r}+A(1 + r)^n-A"


A = the annuity payment , r = the interest rate per time period, and n = the number of time periods.


Substituting values

"S=X×\frac{(1 + \frac{k}{12})^{24} - 1}{\frac{k}{12}}+X(1 + \frac{k}{12})^{24}-X" in this case S =future value X = the annuity payment or periodic rent, k= the interest rate per time period, and 24 = the number of time periods.